# LAMA-UT: Language Agnostic Multilingual ASR Through Orthography Unification and Language-Specific Transliteration

### v3

@@ Introduction
Since the revolution in ASR technologies driven by self-supervised learning (SSL) models (Baevski et al. 2020; Conneau et al. 2020; Hsu et al. 2021), monolingual ASR has achieved superhuman transcription performance, shifting the main focus of recent research towards developing a universal model that spans multiple languages.
@@
Figure 2: Problems derived in single-token IPA recognition. Diacritic ‘:’ indicates phoneme length, which has no explicit phonetic value. Epsilon denotes a blank token in CTC.
@@
Table 2: Comparison between two orthography unification methods. We report PER and CER for seen and unseen languages. Average values are calculated over all 102 seen and 25 unseen languages, respectively. For a fair comparison, all the model sizes are set to 300 million. $\dagger$ denotes results measured by PER, which does not allow for a strict comparison with other results.
@@ Results
We conducted a performance comparison between our universal transcription generator and the existing baseline, wav2vec2.0-phoneme (Xu, Baevski, and Auli 2021). Our universal transcription generator focuses on generating character-level tokens and is measured using Character Error Rate (CER), while the baseline wav2vec2.0-phoneme is measured using Phoneme Error Rate (PER). However, since both metrics are fundamentally used for estimating phonetic symbols, this comparison can be considered meaningful. Following the Table 2, the results show that the proposed method demonstrated significantly better performance across a broader range of languages compared to existing approaches even not utilizing language-specific modules (e.g. n-gram LM). Furthermore, our pipeline demonstrated relatively strong transcription capabilities for unseen languages that were not explicitly included in the training data. In conclusion, transcribing diverse languages based on their pronunciation can produce a universal transcription which is highly effective.
@@
Among the two methods for standardizing orthographic features, Romanization proved to be more effective than IPA. Its ability to represent pronunciation across languages while reducing complexity makes it a superior choice for meaningful results. Romanization balances phonetic accuracy with simplicity, providing better alignment with LLMs and ensuring efficient processing across multilingual ASR tasks. However, since these results are constrained to the first phase, we have constructed the end-to-end performance comparison between IPA-based and Romanization-based LAMA-UT, and the results are shown in the appendices [^2] https://github.com/sanghyang00/LAMA-UT-Appendices.

### v4

@@ Introduction
Since the revolution in ASR technologies driven by self-supervised learning (SSL) models (Baevski et al. 2020; Conneau et al. 2020; Hsu et al. 2021; Chen et al. 2022), monolingual ASR has achieved superhuman transcription performance, shifting the main focus of recent research towards developing a universal model that spans multiple languages.
@@
Figure 2: Problems derived in single-token IPA recognition. Diacritic ‘:’ indicates phoneme length, which has no explicit phonetic value. $\epsilon$ denotes a blank token in CTC.
@@
Table 2: Comparison between two orthography unification methods. We report PER and CER for seen and unseen languages. Average values are calculated over all 102 seen and 25 unseen languages, respectively. For a fair comparison, all the model sizes are set to 300 million. $\dagger$ denotes results measured by PER, which does not allow for a strict comparison with other results. The language codes and their corresponding language names are provided in the appendix.
@@ Results
We conducted a performance comparison between our universal transcription generator and the existing baseline, wav2vec2 phoneme (Xu, Baevski, and Auli 2021). Our universal transcription generator focuses on generating character-level tokens and is measured using Character Error Rate (CER), while the baseline wav2vec2 phoneme is measured using Phoneme Error Rate (PER). However, since both metrics are fundamentally used for estimating phonetic symbols, this comparison can be considered meaningful. Following the Table 2, the results show that the proposed method demonstrated significantly better performance across a broader range of languages compared to existing approaches even not utilizing language-specific modules (e.g. n-gram LM). Furthermore, our pipeline demonstrated relatively strong transcription capabilities for unseen languages that were not explicitly included in the training data. In conclusion, transcribing diverse languages based on their pronunciation can produce a universal transcription which is highly effective.
@@
Among the two methods for standardizing orthographic features, Romanization proved to be more effective than IPA. Its ability to represent pronunciation across languages while reducing complexity makes it a superior choice for meaningful results. Romanization balances phonetic accuracy with simplicity, providing better alignment with LLMs and ensuring efficient processing across multilingual ASR tasks. However, since these results are only constrained to the first phase of the pipeline, we have constructed the full end-to-end performance comparison between IPA-based and Romanization-based LAMA-UT, and the results are shown in the appendix.
@@ Conclusion
Appendix A Appendix
The following appendix sections propose additional experimental results and specifications for LAMA-UT, which can be instrumental in deeper understanding or validating the strengths of the proposed pipeline. The first section of the appendix presents the specifications of the languages we leveraged in the experiments, given the corresponding ISO-639 language code. Consequently, the second section demonstrates the transcription performance of LAMA-UT for each language used in the experiments. Then, the third section suggests the end-to-end transcription performance comparison between IPA-based LAMA-UT and Romanization-based LAMA-UT, further to validate the effectiveness of Romanization in orthography unification. Finally, in the last section, we would like to discuss possible advancements of LAMA-UT and our further research direction.
Appendix B Languages and Language Codes
Seen
Afrikaans (af), Amharic (am), Arabic (ar), Assamese (as), Asturian (ast), Azerbaijani (az), Belarusian (be),
Bulgarian (bg), Bengali (bn), Bosnian (bs), Catalan (ca), Cebuano (ceb), Sorani-Kurdish (ku), Mandarin Chinese (cmn),
Czech (cs), Welsh (cy), Danish (da), German (de), Greek (el), English (en), Spanish (es), Estonian (et), Persian (fa),
Fula ($f$), Finnish (fi), Filipino (fil), French (fr), Irish (ga), Galician (gl), Gujarati (gu), Hausa (ha), Hebrew (he),
Hindi (hi), Croatian (hr),Hungarian (hu), Armenian (hy), Indonesian (id), Igbo (ig), Icelandic (is), Italian (it),
Japanese (ja), Javanese (jv), Georgian (ka), Kamba (kam),Kabuverdianu (kea), Kazakh ($k$), Khmer (km), Kannada (kn),
Korean (ko), Kyrgyz (ky), Luxembourgish (lb), Ganda (lg), Lingala (ln), Lao (lo), Lithuanian (lt), Luo (luw), Latvian (lv),
Maori (mi), Macedonian (mk), Malayalam (ml), Mongolian (mn), Marathi (mr), Malay (ms), Maltese (mt), Burmese (my),
Norwegian (no), Nepali (ne), Dutch (nl), Northern-Sotho (nso), Nyanja (ny), Occitan (oc), Oromo (om), Oriya (or),
Punjabi (pa), Polish (pl), Pashto (ps), Portuguese (pt), Romanian (ro), Russian (ru), Sindhi (sd), Slovak (sk), Slovenian (sl),
Shona (sn), Somali (so), Serbian (sr), Swedish (sv), Swahili (sw), Tamil (ta), Telugu (te), Tajik (tg), Thai (th), Turkish (tr),
Ukrainian (uk), Umbundu (umb), Urdu (ur), Uzbek (uz), Vietnamese (vi), Wolof (wo), Xhosa (xh), Yoruba (yo),
Cantonese Chinese (yue), Zulu (zu)
Unseen
Abkhazian (ab), Albanian (sq), Basaa (bas), Bashkir (ba), Basque (eu), Breton (br), Chuvash (cv), Eastern Mari (mhr),
Erzya (myv), Esperanto (eo), Guarani (gn), Hakha Chin (cnh), Interlingua (ia), Kinyarwanda (rw), Latgalian (ltg),
Norwegian Nynorsk ($n$), Romansh (rm), Tatar ($t$), Toki Pona (tok), Turkmen (tk), Uighur (ug), Upper Sorbian (hsb),
Western Frisian (fy), Western Mari (mrj), Yakut (sah)
Table 8: Specifications of 102 seen languages and 25 unseen languages for experiments. We primarily reported ISO-639-1 codes and provided ISO-639-3 codes when the former were unavailable. Sursilvan and Vallader dialects are both considered Romansh languages.
Table 8 indicates the languages that were leveraged in the experiments of our manuscript. All of the seen languages comprise 102 languages in the FLEURS dataset, and they are used in both training and evaluation of LAMA-UT. Evaluation samples for unseen languages were chosen from the official test split of the CommonVoice 17.0 dataset, which ensured that they possessed an adequate volume of data.
Appendix C Language-Specific Performance of LAMA-UT
Language CER $\downarrow$ Language CER $\downarrow$ Language CER $\downarrow$ Language CER $\downarrow$
Seen Afrikaans 19.3 Ganda 10.6 Lithuanian 8.7 Shona 9.0
Amharic 8.7 Georgian 9.4 Luo 7.4 Sindhi 29.7
Arabic 8.8 German 8.4 Luxembourgish 13.6 Slovak 5.4
Armenian 5.6 Greek 11.5 Macedonian 5.3 Slovenian 9.1
Assamese 12.9 Gujarati 8.9 Malay 8.3 Somali 16.8
Asturian 9.3 Hausa 9.7 Malayalam 7.7 Sorani-Kurdish 11.8
Azerbaijani 10.7 Hebrew 20.5 Maltese 8.4 Spanish 4.7
Belarusian 8.0 Hindi 11.6 Mandarin Chinese 6.5 Swahili 6.4
Bengali 9.8 Hungarian 14.5 Maori 8.7 Swedish 11.8
Bosnian 6.9 Icelandic 18.7 Marathi 11.9 Tajik 5.3
Bulgarian 7.6 Igbo 14.0 Mongolian 14.7 Tamil 11.9
Burmese 19.4 Indonesian 6.0 Nepali 11.2 Telugu 10.8
Cantonese Chinese 16.6 Irish 28.4 Northern-Sotho 13.8 Thai 15.9
Catalan 7.5 Italian 3.7 Norwegian 9.2 Turkish 6.4
Cebuano 7.5 Japanese 33.4 Nyanja 12.4 Ukrainian 11.2
Croatian 6.3 Javanese 7.0 Occitan 17.7 Umbundu 9.5
Czech 11.0 Kabuverdianu 6.8 Oriya 14.3 Urdu 44.0
Danish 15.6 Kamba 10.9 Oromo 20.2 Uzbek 16.3
Dutch 11.3 Kannada 7.4 Pashto 20.0 Vietnamese 13.0
English 13.1 Kazakh 5.1 Persian 6.9 Welsh 12.8
Estonian 4.4 Khmer 23.6 Polish 10.9 Wolof 14.9
Filipino 4.8 Korean 9.6 Portuguese 9.0 Xhosa 10.4
Finnish 4.9 Kyrgyz 5.2 Punjabi 20.3 Yoruba 10.9
French 12.3 Lao 17.8 Romanian 11.8 Zulu 9.5
Fula 14.7 Latvian 7.0 Russian 9.0
Galician 6.7 Lingala 7.7 Serbian 8.9
Unseen Abkhazian 45.8 Eastern Mari 30.9 Latgalian 28.8 Upper Sorbian 35.3
Albanian 34.8 Erzya 29.9 Norwegian Nynorsk 27.1 Western Frisian 33.8
Basa 33.4 Esperanto 22.1 Romansh 31.9 Western Mari 36.7
Bashkir 33.5 Guarani 33.5 Tatar 31.8 Yakut 36.3
Basque 33.0 Hakha Chin 42.9 Toki Pona 28.7
Breton 49.1 Interlingua 17.3 Turkmen 41.3
Chuvash 37.4 Kinyarwanda 37.2 Uighur 27.1
Table 9: Performance of Romanization-based universal transcription generator in 102 seen and 25 unseen languages. Results are reported in CER, which calculates the character-level difference between ground truth Romanized transcription and predicted Romanized transcription.
Language CER $\downarrow$ Language CER $\downarrow$ Language CER $\downarrow$ Language CER $\downarrow$
Seen Afrikaans 18.3 Ganda 10.4 Luo 7.3 Sindhi 29.1
Amharic 47.4 Georgian 23.6 Luxembourgish 16.5 Slovak 3.8
Arabic 13.9 German 5.5 Macedonian 5.5 Slovenian 9.4
Armenian 22.4 Greek 15.6 Malay 6.9 Somali 15.4
Assamese 18.9 Gujarati 11.9 Malayalam 24.1 Sorani-Kurdish 26.4
Asturian 9.8 Hausa 10.2 Maltese 9.8 Spanish 2.8
Azerbaijani 11.9 Hebrew 35.2 Mandarin Chinese 36.0 Swahili 5.3
Belarusian 11.3 Hindi 8.3 Maori 12.0 Swedish 9.5
Bengali 13.8 Hungarian 16.7 Marathi 14.0 Tajik 8.9
Bosnian 5.9 Icelandic 18.1 Mongolian 21.4 Tamil 19.6
Bulgarian 8.2 Igbo 17.9 Nepali 13.2 Telugu 15.5
Burmese 52.6 Indonesian 4.3 Northern-Sotho 15.1 Thai 45.3
Cantonese Chinese 61.6 Irish 34.8 Norwegian 7.0 Turkish 5.1
Catalan 6.3 Italian 2.0 Nyanja 12.1 Ukrainian 9.7
Cebuano 7.4 Javanese 6.6 Occitan 19.9 Umbundu 10.5
Croatian 5.2 Kabuverdianu 8.3 Oriya 18.3 Urdu 14.9
Czech 9.9 Kamba 14.6 Oromo 19.8 Uzbek 14.5
Danish 12.9 Kannada 13.9 Pashto 32.7 Vietnamese 19.0
Dutch 8.7 Kazakh 6.7 Persian 11.9 Welsh 12.3
English 8.2 Khmer 66.3 Polish 8.6 Wolof 18.0
Estonian 5.7 Korean 18.8 Portuguese 7.9 Xhosa 10.5
Filipino 4.4 Kyrgyz 8.8 Punjabi 19.5 Yoruba 25.5
Finnish 4.7 Lao 55.0 Romanian 11.0 Zulu 9.6
French 8.2 Latvian 7.9 Russian 6.0
Fula 16.6 Lingala 7.8 Serbian 5.7
Galician 6.0 Lithuanian 8.6 Shona 8.6
Unseen Abkhazian 55.9 Eastern Mari 35.3 Latgalian 34.7 Upper Sorbian 35.2
Albanian 35.4 Erzya 35.7 Norwegian Nynorsk 23.9 Western Frisian 29.6
Basa 42.1 Esperanto 19.7 Romansh 30.0 Western Mari 41.4
Bashkir 29.7 Guarani 40.0 Tatar 24.2 Yakut 38.2
Basque 32.8 Hakha Chin 42.8 Toki Pona 29.0
Breton 49.9 Interlingua 15.2 Turkmen 43.1
Chuvash 44.3 Kinyarwanda 36.7 Uighur 24.5
Table 10: Performance of LAMA-UT in 101 seen and 25 unseen languages. We leveraged a Romanization-based universal transcription generator and GPT-4o-mini model with a few-shot prompting as a universal converter. All of the results are reported in CER, which calculates the character-level difference between ground truth language-specific transcription and predicted language-specific transcription. Japanese was excluded from the results due to the ambiguity in evaluation criteria arising from its mixed use of Hiragana, Katakana, and Kanji.
In this section, we present a detailed performance analysis for each language, comparing transcription performance from the universal transcription generation phase to that after passing through the universal converter. Additionally, we analyze these results further to explore the correlation between orthographic characteristics and transcription performance. First of all, Table  9 suggests the language-specific transcription performance of the Romanization-based universal transcription generator, and Table 10 presents the language-specific end-to-end (E2E) language-specific transcription performance of LAMA-UT. In the case of the universal transcription generator, CERs were consistently similar across most languages with minimal variation. However, in the end-to-end pipeline with language-specific transliteration, notably higher error rates were observed for certain languages, and the languages highlighted in bold represent the top 10 languages with the most significant increases in error rate observed in the end-to-end pipeline compared to the CER values during the universal transcription phase. Interestingly, the 10 languages that exhibited the most significant performance degradation after passing through the universal converter were all non-Latin script languages, with the majority being languages that do not employ spacing in their orthography. This suggests that when a pre-trained LLM is utilized as a universal converter, languages with non-Latin orthography which inherently exhibit different structural characteristics compared to Latin-based languages, are more prone to error propagation.
Appendix D E2E Comparison on Unification Methods
Universal
Converter
IPA Roman
Upper Bound Few Shot Upper Bound Few Shot
CER $\downarrow$ WER $\downarrow$ CER $\downarrow$ WER $\downarrow$ CER $\downarrow$ WER $\downarrow$ CER $\downarrow$ WER $\downarrow$
Seen LLaMA-8B 51.4 82.4 56.5 80.8 33.0 44.4 38.1 64.7
LLaMA-70B 40.0 64.6 33.8 67.8 15.6 34.6 23.1 56.2
GPT-4o-mini 27.3 49.2 33.0 70.6 10.0 28.6 19.9 49.4
Unseen LLaMA-8B 52.0 93.0 43.4 93.3 32.7 60.2 41.6 94.4
LLaMA-70B 42.5 89.0 38.1 94.6 17.0 57.5 34.7 93.9
GPT-4o-mini 40.4 78.2 35.2 84.1 12.3 42.1 29.7 83.0
Table 11: An end-to-end (E2E) comparison between IPA and Romanization. Since Phonemizer (IPA transliterator) supports a smaller amount of language compared to Uroman (Romanization transliterator), all of the metrics were calculated and averaged within the common languages which are both available in the IPA-based and the Romanization-based version of the universal transcription generator.
In this section, we would like to further clarify the effectiveness of utilizing Romanization as an intermediate representation. Since the experimental results from Table 2 were limited to the universal transcription generation phase, we supplement the previous results with the full performance comparison between IPA-based and Romanization-based end-to-end architecture of LAMA-UT, both in upper bound assessment and transcription performance. Table 11 proposes the end-to-end performance comparison between IPA-based LAMA-UT and Romanization-based LAMA-UT. The only difference between the two models is the orthography unification method utilized in the universal transcription generator. (e.g., the IPA-based model utilizes IPA as an intermediate representation, while the Romanization-based one leverages Romanized transcription as an intermediate representation. In the results, Romanization-based LAMA-UT consistently outperforms IPA-based LAMA-UT as a substantial difference in both CER and WER. These results strongly demonstrate the superiority of Romanization over IPA when leveraging LLMs as a universal converter, since LLMs are primarily trained with Latin scripts and optimized tokenization strategy for them. Furthermore, there are some inconsistencies in the results of the IPA-based LAMA-UT, where passing predicted IPA transcriptions along with a few examples to the universal converter yielded better performance than passing ground truth IPA transcriptions without examples. This is presumably because the LLMs did not frequently encounter IPA-driven tokens during its training process.
Appendix E Discussion and Future Work
LAMA-UT showed comparable or better performance compared to previous works even without language-specific modules (e.g., adapters, lexicons, n-gram LMs), while achieving efficient training with a significantly reduced dataset size. However, there is still room for further improvement in both the universal transcription generator and the universal converter. First, the transcription performance of the universal transcription generator can be improved. For instance, the universal transcription generator of LAMA-UT can leverage additional linguistic information (e.g., embedding from a pre-trained language classifier) to further enhance the transcription quality of the first phase. Secondly, our pipeline shows relatively lower performance for languages with distinct linguistic structures, like Korean, and those with additional features (e.g., tones), such as Chinese, in the language-specific transliteration phase. Since our universal converter is replaceable, this issue will naturally be resolved in line with the development of LLMs. Finally, the utilization of prompt learning techniques (Li and Liang 2021; Liu et al. 2022; Gu et al. 2022) might improve transliteration performance. We plan to address these aspects in future research.